OpenStudy (anonymous):

show me how to calculate -sin(3pi/4)

OpenStudy (anonymous):

you want to get the reference angle, that is the first quadrant angle that has the same value, maybe varying only by sign

OpenStudy (anonymous):

we know the sign is positive in the 2nd quadrant and in the first quadrant

OpenStudy (anonymous):

of sine, that is

OpenStudy (anonymous):

so this question is the exact same thing as -sin(3pi/4-pi/2)

OpenStudy (anonymous):

which is -sin(pi/4)

OpenStudy (anonymous):

think of the unit circle with radius 1, what would the y-value need to be with respect to the radius for theta to be pi/4?

OpenStudy (mertsj):

 So \[-\sin\frac{3\pi}{4}=-\sin\frac{\pi}{4}=-\frac{\sqrt{2}}{2}\]